OpenStudy (anonymous):

Gross domestic product (GDP) is a measure of the economic output of a country. The GDP of the United States was about 1.2 x 10^13 dollars in 2005. This is about 3 times the GDP in 1985. What was the U.S. GDP in 1985? 4 x 10^2 4 x 10^3 4 x 10^12 4 x 10^13

OpenStudy (jamesj):

\[ GDP_{1985} = \frac{1}{3} GDP_{2005} \] \[ = \frac{1}{3} (1.2 \times 10^{13}) \] \[ = \frac{1.2}{3} \times 10^{13} \] \[ = 0.4 \times 10^{13} \] Now which of the solutions is that equal to?

OpenStudy (anonymous):

the bottom one thanks

OpenStudy (jamesj):

No, that's not right.